OpenStudy (anonymous):

49c^5-7c^2/-7c explain.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so I'll split things up for you \[\frac{49c^4 \times c}{-7c} - \frac{7c \times c}{-7c}\] can you seed anything in th denominator that cancels with the numerator..?

OpenStudy (anonymous):

7c?

OpenStudy (anonymous):

or 7x7 is 49 idk

OpenStudy (campbell_st):

great 49c/-7c = -7 and -7c/-7c = c so \[\frac{49c^5 - 7c^2}{-7c} = -7c^4 + c\]

OpenStudy (anonymous):

oh okay, and i have a couple more i need help with

OpenStudy (campbell_st):

there are lots of people who can help... just post them in the normal way

OpenStudy (anonymous):

yeah and then i have to wait like an hour before someone actually looks at it

OpenStudy (anonymous):

and i dont want to wait alllllllllll day just to wait for help on a couple problems lol

OpenStudy (anonymous):

-3a^4-3a/-3a

OpenStudy (anonymous):

can you outline it like the one i just did cause its hard to solve it this way

OpenStudy (anonymous):

@NoelGreco can you help?